# AI-Assisted Object Condensation Clustering for Calorimeter Shower Reconstruction at CLAS12

### v1

@@ 7 Conclusion
This paper presents a novel AI approach to ECal hit clustering at CLAS12 using object condensation. Our model uses both GravNet layers for local message-passing and a Transformer encoder for long-range message-passing. The trained network was integrated into the existing CLAS12 reconstruction pipeline, where it was shown to outperform previous methods in producing reliable neutron and photon clusters.
@@ Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under contract DE-SC0024505. This research was supported by the U.S. Department of Energy (DOE) and the National Science Foundation (NSF). We extend our gratitude to the DOE and NSF for their financial support, which made this work possible.

### v3

@@ 7 Conclusion
This paper presents a novel AI approach to ECal hit clustering at CLAS12 using object condensation. Our model uses both GravNet layers for local message-passing and a Transformer encoder for long-range message-passing. The trained network was integrated into the existing CLAS12 reconstruction pipeline, where it was shown to outperform previous methods in producing reliable neutron and photon clusters by a trustworthiness metric. This improved trustworthiness can relax the need for overly conservative selection cuts in a range of exclusive measurements and opens the door for previously inaccessible neutron semi-inclusive DIS measurements.
@@ Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under contracts DE-SC0024505 and DE-AC05-06OR23177. This research was supported by the National Science Foundation (NSF) under grant agreement DGE-2139754. We extend our gratitude to the DOE and NSF for their financial support, which made this work possible. We also thank the CLAS collaboration for providing the reconstruction framework and simulation environment for this project.